OpenStudy (anonymous):

GIVING MEDALS AND FANING PEOPLE!!! Write the given equation in slope-intercept form, y = mx + b, by solving for y. 4x – 2y = 10

OpenStudy (anonymous):

4x - 2y = 10 The first thing you want to do is isolate y on one side. So, subtract 4x from both sides. 4x - 2y = 10 -4x - 4x -2y = -4x + 10 Divide each side by -2. -2y/-2 = -4x/-2 +10/-2 y = 2x - 5 Is this understandable? (:

OpenStudy (anonymous):

yes.. omg thank you so much :)

OpenStudy (anonymous):

You're welcome. (: